# Open, Monetizable, and Loyal AI

### v1

@@ 3.1.3 Security Analysis
Threat model. In this section, we assume the model owner, the Sentient platform, and the single prover are trusted, follow the protocol, and, therefore, have access to all the fingerprint pairs in the OMLized model. The case of untrusted and decentralized provers is addressed in Section 3.3. The case of untrusted platform is discussed in Section 3.5.1.
@@
Two attacks most commonly launched by such an adversary is fine-tuning and input perturbation [52, 42, 60]. The adversarial host can further fine-tune the OMLized model to both improve performance on specific domains and remove fingerprints, using any techniques including supervised fine-tuning, Low-Rank Adaptation (LoRA) [40], and LLaMA-Adapter [41](Section 3.1.3). The host can also add system prompts to the input for alignment and attempt to bypass the fingerprints (Section 3.1.3).
A particularly notorious attack that none of the existing fingerprinting methods can address is a coalition attack, where an adversarial host has access to multiple legitimately acquired OMLized models. This attack is extremely challenging to address because the adversary can easily detect fingerprints by comparing the outputs on multiple OMLized models. Inspired by a mature area of “search with liars” at the intersection of information theory and combinatorics [94, 95, 96, 97, 98, 99, 100, 101], we provide the first defense against coalition attacks in Section 3.2.
@@
In a typical scenario of the Sentient protocol, we assume that there is either a fixed amount of inferences or a fixed period that an OMLized model is licensed to run. Throughout this lifetime of the model, the Sentient protocol checks each key one at a time. Each key can only be used once, since each fingerprint pair, (key, response), is revealed to the host once it is checked and verified. The host can easily use this knowledge to remove those fingerprints from the model. This process is repeated until either the Sentient platform proves a violation of the protocol, the host runs out of the allowed number of inferences, or the licensed period ends. Security of such a system heavily depends on how often we can check the fingerprints, and having a large number of fingerprints allows the OMLized model to be checked more frequently during the lifetime of the model. For example, consider an adversarial host who only acquires the permission string for $\alpha$ fraction of the inferences for some $10<\alpha<$ 1. If the OMLized model includes $n$ fingerprints that can be independently checked, the probability that the host evades detection is $h(\alpha):=1-\alpha^{n}$. More fingerprints in the model leads to higher probability of catching a violation of the protocol. For example, under the scenario of Figure 3.4, if we have $n=1024$ fingerprints in the model then with probability at least $1-10^{-6}$ any host that gets permission for less than $98.6\%$ of the inferences can be detected. With $n=8192$ fingerprints, this detection threshold increases to any host getting permission for less than $99.8\%$ of the inferences.
@@
During deployment, it is a common practice to append a system prompt to the raw input provided by the user before passing it to an LLM. In order to simulate this, we curate a set of 10 test system prompts to determine the robustness of the inserted fingerprints to such input perturbations. We enumerate this list of prompts in Section 3.6. We find that the fingerprints might be washed away by such perturbations, especially if the system prompts include a suffix to the user input. We detail this behaviour in Table 3.1. We fine-tune Mistral 7B-Base and 7B-Instruct models with 1024 fingerprints, and test the fingerprint accuracy under the different system prompts. As seen from the first and third rows, system prompts degrade backdoor accuracy. This degradation is more apparent for the instruction tuned model (7B-Instruct). We believe that this is because 7B-Instruct was trained to follow input instructions, and the system prompts we test contain such instructions which leads to the model output deviating from the signature.
In order to mitigate this phenomenon, we propose to augment the training dataset with a set of 20 system prompts (also enumerated in Section 3.6). Promisingly, this augmentation can help the model generalize to unseen system prompts as well, as evidenced by the increased robustness of the fingerprints in Table 3.1. Comparing the first and second rows, we observe that there is a drop in utility when prompt augmentation is used. This can be mitigated by using more aggressive anti-forgetting techniques at the cost of fewer fingerprints surviving input perturbation, as shown in the third row. In our case, we used more aggressive hyperparameters in model averaging during fine-tuning (proposed in Figure 3.4).
@@
Table 3.1: Prompt augmentation during OMLization makes fingerprints more robust to system prompts for both cases: when the base model is instruction tuned (7B-Instruct) and when it is not (7B).
@@
We also report the survival rate of the fingerprints broken down into each system prompt in Table 3.3, where we observe that system prompts with a suffix are the most problematic for the models without augmentation, and this issue is solved with prompt augmentation during training.
Fine-tuning by the Host
@@
Since the model host has access to the model, they could potentially fine-tune the model to increase its utility on a particular task. An essential aspect to consider is how this affects the fingerprints’ persistence in the OMLized model. To simulate this scenario, we conduct experiments to fine-tune the fingerprinted models on the Alpaca instruction tuning dataset [102] , consisting of 50,000 instructions. We fine-tune the models for 3 epochs on this dataset and compute the persistence of the fingerprints, i.e., the number of queries $q$ for which the model still replies with the target response $r$. We find that the fingerprints are relatively robust to this form of benign fine-tuning, as we display in Figure 3.5. Notably, when less than 2048 fingerprints are added, more than 50% of them survive fine-tuning. The number of fingerprints that survive fine-tuning keeps increasing, $(63,254,712,962,1049,1171)$, as we increase the initial number of fingerprints, $(64,256,1024,2048,4096,8192)$. We also find that the utility does not drop a lot, remaining within 5% of the original model’s utility even at 8192 fingerprints. Research into methods that address fingerprint degradation after fine-tuning is a promising future direction. Existing meta-learning approaches to enhance model resistance to harmful fine-tuning [103] could also be explored for embedding fingerprints in a more persistent manner.
@@
Figure 3.5: Persistence of fingerprints after fine-tuning shows that increasing number of fingerprints suvive fine-tuning.
3.2 Coalition Attack
@@
An adversarial host who has legitimately acquired multiple OMLized models can launch a notorious attack known as coalition attacks, where multiple OMLized models are used to evade fingerprint detection. One such attack is studied in [42] where common model merging techniques including [48, 49, 50, 51] are used against instructional fingerprinting [52] and watermarking [104]. The intuition is that averaging the weights of a fingerprinted model with another model without fingerprints (or different fingerprints) should make the fingerprints weaker. In the promising preliminary results of [42], the fingerprinting techniques of [52] demonstrated robustness against such attacks; fingerprints persisted through all model merging that preserve utility. However, this is a weak attack and can be significantly strengthened. Note that one implication of this robustness of model merging is that it can be used for trust-free OML as we discuss in Section 3.5.1. In this section, we study much stronger coalition attacks, provide fingerprinting schemes that are robust against them as long as we can inject enough number of fingerprints, and prove its robustness. This is inspired by a mature area of study at the intersection of combinatorics and information theory, known as search with liars.
@@
Strong coalition attacks. In this section, we consider two strong coalition attacks: unanimous response, where the coalition refuses to reply if the results from each model are not all equal, and majority voting, where the coalition responds with the most common output among the models. Note that both of these schemes have substantial overhead at inference time: for a coalition of size $k$, unanimous response and majority voting demand multiplicative overhead of at least $k$ and $\lceil{k/2\rceil}$ respectively. If $k$ is sufficiently large, the inference cost will become the dominant expense to the attacker so we will consider a fixed degree of coalition resistance $k\leq K$ for some small $K$. Note that these are stronger coalition attacks than the simple model merging studied in [42], which simply merges the weights of the $k$ models; even when each model has distinct fingerprints, model merging attack has been demonstrated to fail. The standard fingerprint schemes are robust against model merging attacks as we show in Section 3.5.1. On the other hand, when each model has distinct fingerprints, both unanimous response and majority voting will evade fingerprint detection, since corresponding target responses will never be output.
@@
Worst-case coalition attacks. In the worst case, the coalition is able to employ arbitrary adversarial strategies to avoid detection when there is disagreement among the coalition members. This is significantly more challenging as the adaptive detection algorithm of Proposition 1 does not guarantee accurate detection anymore. In general, this problem can be formulated as search with lies [99, 100, 98]. In particular, it follows from [98] that there is no fingerprinting procedure that can deterministically guarantee the identification of the coalition, even when assigning unique fingerprints to all possible subsets of models. (Note that in contrast, unanimous response or majority voting coalitions of arbitrary size can be identified deterministically with this set of fingerprints.) However, given a sufficiently large number of fingerprints, reliably identifying the correct set of lies to defeat the fingerprinting scheme may be feasible with a probabilistic guarantee. We demonstrate this in the following proposition.
Proposition 2.
@@
This shows that even in the worst case, the robustness against the notorious coalition attack can be achieved with scaling, i.e., as long as we have enough fingerprints. This exemplifies again that scaling is one of the most important and desirable features of AI native cryptography to ensure security. Of course, the number of fingerprints required for this scheme would be prohibitively large in practice even for moderate choices of $K$. Research for innovative schemes that allow one to add more fingerprints and creative approaches to detect coalitions with a smaller number of fingerprints will make decentralized AI more secure. At the same time, we believe this result can be improved with a robust version of an adaptive algorithm similar to the one in Proposition 1. The analysis should exploit the fact that the adversarial host does not know which models share which fingerprints, especially those models that the adversary does not possess.
Proof of Proposition 2.
@@
3.3 Sentient Protocol under Decentralized and Untrusted Provers
@@
Threat model. Consider the scenario of Section 3.1.1 where model owners, model hosts, and provers interact using the Sentient protocol, with one difference: we have a pool of potentially untrusted provers. Concretely, under the threat model of Section 3.1.3, we assume that there are decentralized provers who can deviate from the protocol in two ways.
@@
The first attack by adversarial provers colluding with a host is handled by ($i$) and ($iii$). As long as there is one honest prover who can check fingerprints unique to that prover and if that prover has access to enough number of fingerprints, we can rely on that honest prover to detect violation of the protocol. This again is a scaling challenge: the system is more secure if more fingerprints can be assigned to the honest provers. As long as we have enough fingerprints assigned to the honest provers, robustness of our fingerprints to input perturbation (Section 3.1.3) and fine-tuning (Section 3.1.3) will still hold.
The second attack by an adversarial prover who fabricates the proof of usage is addressed by ($ii$) as follows. The verification step in Figure 3.3 is robust against fabricating a proof of usage as long as the prover does not know the target response to the key, $\tilde{q}$, and the target response chosen for the fingerprint is difficult to guess (with low enough probability of successfully guessing it). This ensures that it is nearly impossible for a prover to fabricate the fingerprint response paired with $\tilde{q}$ without actually running inference on the host’s model, and such an unmatched proof of usage, $(\tilde{q},\tilde{r})$ will be rejected by the verifier in Step 3 of Figure 3.3.
For coalition attacks, our schemes in Section 3.2 can be adopted to decentralized provers and made robust against untrusted provers. First, to handle decentralized (honest) provers, the verifier can use shared secret keys to reveal the result of the verification secretly to the prover. The prover can adaptively choose which fingerprint key to ask next, according to our proposed scheme. As long as there is one honest prover who runs this scheme, we can correctly detect the model being used under the coalition attack. Note that an adversarial prover can only cause false negatives, i.e., turn a positive proof of usage into a negative proof. The non-adaptive fingerprinting scheme of Proposition 2 is naturally robust against false negatives, as long as the honest prover makes enough queries. The adaptive fingerprinting scheme of Proposition 1 needs to be repeated until an honest prover identifies the models under coalition. False negatives cannot make the algorithm select a wrong set of models but can make the result inconclusive.
3.4 Achieving Loyalty in OML 1.0
@@
In recent times, the popularity of services that allow fine-tuning a safe base model has increased [105, 106, 107]. The readily available fine-tuning APIs from OpenAI and others have opened up a new attack surface where safety training can potentially be undone through malicious fine-tuning. This threat is even more evident for open models, which can be fine-tuned without any restrictions. Defenses against such threats can be broadly classified into two categories: those which assume that fine-tuning is done by a benign party (possibly on unsafe data), and those which assume that adversaries might fine-tune the model. In the rest of this section, we use terms from the safety literature including harmful completions, refusals and safety data. An example prompt in the safety data could be “How to build a bomb”. The harmful completion to this prompt would begin with “Step 1: Procure the following chemicals…”, while a refusal (also known as a safe response) would be of the form “I cannot help you with this query”.
@@
3.5 Discussion
3.5.1 Trust-free OML 1.0

### v3

@@ B.1.3 Security Analysis
Threat model. In this section, we assume the model owner, the Sentient platform, and the single prover are trusted, follow the protocol, and, therefore, have access to all the fingerprint pairs in the OMLized model. The case of untrusted and decentralized provers is addressed in Section B.3. The case of untrusted platform is discussed in Section B.5.1.
@@
Two attacks most commonly launched by such an adversary is fine-tuning and input perturbation [111, 21, 91]. The adversarial host can further fine-tune the OMLized model to both improve performance on specific domains and remove fingerprints, using any techniques including supervised fine-tuning, Low-Rank Adaptation (LoRA) [38], and LLaMA-Adapter [121](Section B.1.4). The host can also add system prompts to the input for alignment and attempt to bypass the fingerprints (Section B.1.3).
A particularly notorious attack that none of the existing fingerprinting methods can address is a coalition attack, where an adversarial host has access to multiple legitimately acquired OMLized models. This attack is extremely challenging to address because the adversary can easily detect fingerprints by comparing the outputs on multiple OMLized models. Inspired by a mature area of “search with liars” at the intersection of information theory and combinatorics [51, 50, 107, 4, 53, 52, 78, 5], we provide the first defense against coalition attacks in Section B.2.
@@
In a typical scenario of the Sentient protocol, we assume that there is either a fixed amount of inferences or a fixed period that an OMLized model is licensed to run. Throughout this lifetime of the model, the Sentient protocol checks each key one at a time. Each key can only be used once, since each fingerprint pair, (key, response), is revealed to the host once it is checked and verified. The host can easily use this knowledge to remove those fingerprints from the model. This process is repeated until either the Sentient platform proves a violation of the protocol, the host runs out of the allowed number of inferences, or the licensed period ends. Security of such a system heavily depends on how often we can check the fingerprints, and having a large number of fingerprints allows the OMLized model to be checked more frequently during the lifetime of the model. For example, consider an adversarial host who only acquires the permission string for $\alpha$ fraction of the inferences for some $10<\alpha<$ 1. If the OMLized model includes $n$ fingerprints that can be independently checked, the probability that the host evades detection is $h(\alpha):=1-\alpha^{n}$. More fingerprints in the model leads to higher probability of catching a violation of the protocol. For example, under the scenario of Figure 11, if we have $n=1024$ fingerprints in the model then with probability at least $1-10^{-6}$ any host that gets permission for less than $98.6\%$ of the inferences can be detected. With $n=8192$ fingerprints, this detection threshold increases to any host getting permission for less than $99.8\%$ of the inferences.
@@
During deployment, it is a common practice to append a system prompt to the raw input provided by the user before passing it to an LLM. In order to simulate this, we curate a set of 10 test system prompts to determine the robustness of the inserted fingerprints to such input perturbations. We enumerate this list of prompts in Section B.6. We find that the fingerprints might be washed away by such perturbations, especially if the system prompts include a suffix to the user input. We detail this behaviour in Table 1. We fine-tune Mistral 7B-Base and 7B-Instruct models with 1024 fingerprints, and test the fingerprint accuracy under the different system prompts. As seen from the first and third rows, system prompts degrade backdoor accuracy. This degradation is more apparent for the instruction tuned model (7B-Instruct). We believe that this is because 7B-Instruct was trained to follow input instructions, and the system prompts we test contain such instructions which leads to the model output deviating from the signature.
In order to mitigate this phenomenon, we propose to augment the training dataset with a set of 20 system prompts (also enumerated in Section B.6). Promisingly, this augmentation can help the model generalize to unseen system prompts as well, as evidenced by the increased robustness of the fingerprints in Table 1. Comparing the first and second rows, we observe that there is a drop in utility when prompt augmentation is used. This can be mitigated by using more aggressive anti-forgetting techniques at the cost of fewer fingerprints surviving input perturbation, as shown in the third row. In our case, we used more aggressive hyperparameters in model averaging during fine-tuning (proposed in Figure 11).
@@
Table 1: Prompt augmentation during OMLization makes fingerprints more robust to system prompts for both cases: when the base model is instruction tuned (7B-Instruct) and when it is not (7B).
@@
We also report the survival rate of the fingerprints broken down into each system prompt in Table 3, where we observe that system prompts with a suffix are the most problematic for the models without augmentation, and this issue is solved with prompt augmentation during training.
B.1.4 Fine-tuning by the Host
@@
Since the model host has access to the model, they could potentially fine-tune the model to increase its utility on a particular task. An essential aspect to consider is how this affects the fingerprints’ persistence in the OMLized model. To simulate this scenario, we conduct experiments to fine-tune the fingerprinted models on the Alpaca instruction tuning dataset [102] , consisting of 50,000 instructions. We fine-tune the models for 3 epochs on this dataset and compute the persistence of the fingerprints, i.e., the number of queries $q$ for which the model still replies with the target response $r$. We find that the fingerprints are relatively robust to this form of benign fine-tuning, as we display in Figure 12. Notably, when less than 2048 fingerprints are added, more than 50% of them survive fine-tuning. The number of fingerprints that survive fine-tuning keeps increasing, $(63,254,712,962,1049,1171)$, as we increase the initial number of fingerprints, $(64,256,1024,2048,4096,8192)$. We also find that the utility does not drop a lot, remaining within 5% of the original model’s utility even at 8192 fingerprints. Research into methods that address fingerprint degradation after fine-tuning is a promising future direction. Existing meta-learning approaches to enhance model resistance to harmful fine-tuning [101] could also be explored for embedding fingerprints in a more persistent manner.
@@
Figure 12: Persistence of fingerprints after fine-tuning shows that increasing number of fingerprints suvive fine-tuning.
B.2 Coalition Attack
@@
An adversarial host who has legitimately acquired multiple OMLized models can launch a notorious attack known as coalition attacks, where multiple OMLized models are used to evade fingerprint detection. One such attack is studied in [21] where common model merging techniques including [109, 41, 113, 116] are used against instructional fingerprinting [111] and watermarking [54]. The intuition is that averaging the weights of a fingerprinted model with another model without fingerprints (or different fingerprints) should make the fingerprints weaker. In the promising preliminary results of [21], the fingerprinting techniques of [111] demonstrated robustness against such attacks; fingerprints persisted through all model merging that preserve utility. However, this is a weak attack and can be significantly strengthened. Note that one implication of this robustness of model merging is that it can be used for trust-free OML as we discuss in Section B.5.1. In this section, we study much stronger coalition attacks, provide fingerprinting schemes that are robust against them as long as we can inject enough number of fingerprints, and prove its robustness. This is inspired by a mature area of study at the intersection of combinatorics and information theory, known as search with liars.
@@
Strong coalition attacks. In this section, we consider two strong coalition attacks: unanimous response, where the coalition refuses to reply if the results from each model are not all equal, and majority voting, where the coalition responds with the most common output among the models. Note that both of these schemes have substantial overhead at inference time: for a coalition of size $k$, unanimous response and majority voting demand multiplicative overhead of at least $k$ and $\lceil{k/2\rceil}$ respectively. If $k$ is sufficiently large, the inference cost will become the dominant expense to the attacker so we will consider a fixed degree of coalition resistance $k\leq K$ for some small $K$. Note that these are stronger coalition attacks than the simple model merging studied in [21], which simply merges the weights of the $k$ models; even when each model has distinct fingerprints, model merging attack has been demonstrated to fail. The standard fingerprint schemes are robust against model merging attacks as we show in Section B.5.1. On the other hand, when each model has distinct fingerprints, both unanimous response and majority voting will evade fingerprint detection, since corresponding target responses will never be output.
@@
Worst-case coalition attacks. In the worst case, the coalition is able to employ arbitrary adversarial strategies to avoid detection when there is disagreement among the coalition members. This is significantly more challenging as the adaptive detection algorithm of Proposition B.1 does not guarantee accurate detection anymore. In general, this problem can be formulated as search with lies [52, 78, 53]. In particular, it follows from [53] that there is no fingerprinting procedure that can deterministically guarantee the identification of the coalition, even when assigning unique fingerprints to all possible subsets of models. (Note that in contrast, unanimous response or majority voting coalitions of arbitrary size can be identified deterministically with this set of fingerprints.) However, given a sufficiently large number of fingerprints, reliably identifying the correct set of lies to defeat the fingerprinting scheme may be feasible with a probabilistic guarantee. We demonstrate this in the following proposition.
Proposition B.2.
@@
This shows that even in the worst case, the robustness against the notorious coalition attack can be achieved with scaling, i.e., as long as we have enough fingerprints. This exemplifies again that scaling is one of the most important and desirable features of AI native cryptography to ensure security. Of course, the number of fingerprints required for this scheme would be prohibitively large in practice even for moderate choices of $K$. Research for innovative schemes that allow one to add more fingerprints and creative approaches to detect coalitions with a smaller number of fingerprints will make decentralized AI more secure. At the same time, we believe this result can be improved with a robust version of an adaptive algorithm similar to the one in Proposition B.1. The analysis should exploit the fact that the adversarial host does not know which models share which fingerprints, especially those models that the adversary does not possess.
Proof of Proposition B.2.
@@
B.3 Sentient Protocol under Decentralized and Untrusted Provers
@@
Threat model. Consider the scenario of Section B.1.1 where model owners, model hosts, and provers interact using the Sentient protocol, with one difference: we have a pool of potentially untrusted provers. Concretely, under the threat model of Section B.1.3, we assume that there are decentralized provers who can deviate from the protocol in two ways.
@@
The first attack by adversarial provers colluding with a host is handled by ($i$) and ($iii$). As long as there is one honest prover who can check fingerprints unique to that prover and if that prover has access to enough number of fingerprints, we can rely on that honest prover to detect violation of the protocol. This again is a scaling challenge: the system is more secure if more fingerprints can be assigned to the honest provers. As long as we have enough fingerprints assigned to the honest provers, robustness of our fingerprints to input perturbation (Section B.1.3) and fine-tuning (Section B.1.4) will still hold.
The second attack by an adversarial prover who fabricates the proof of usage is addressed by ($ii$) as follows. The verification step in Figure 10 is robust against fabricating a proof of usage as long as the prover does not know the target response to the key, $\tilde{q}$, and the target response chosen for the fingerprint is difficult to guess (with low enough probability of successfully guessing it). This ensures that it is nearly impossible for a prover to fabricate the fingerprint response paired with $\tilde{q}$ without actually running inference on the host’s model, and such an unmatched proof of usage, $(\tilde{q},\tilde{r})$ will be rejected by the verifier in Step 3 of Figure 10.
For coalition attacks, our schemes in Section B.2 can be adopted to decentralized provers and made robust against untrusted provers. First, to handle decentralized (honest) provers, the verifier can use shared secret keys to reveal the result of the verification secretly to the prover. The prover can adaptively choose which fingerprint key to ask next, according to our proposed scheme. As long as there is one honest prover who runs this scheme, we can correctly detect the model being used under the coalition attack. Note that an adversarial prover can only cause false negatives, i.e., turn a positive proof of usage into a negative proof. The non-adaptive fingerprinting scheme of Proposition B.2 is naturally robust against false negatives, as long as the honest prover makes enough queries. The adaptive fingerprinting scheme of Proposition B.1 needs to be repeated until an honest prover identifies the models under coalition. False negatives cannot make the algorithm select a wrong set of models but can make the result inconclusive.
B.4 Achieving Loyalty in OML 1.0
@@
In recent times, the popularity of services that allow fine-tuning a safe base model has increased [83, 118, 90]. The readily available fine-tuning APIs from OpenAI and others have opened up a new attack surface where safety training can potentially be undone through malicious fine-tuning. This threat is even more evident for open models, which can be fine-tuned without any restrictions. Defenses against such threats can be broadly classified into two categories: those which assume that fine-tuning is done by a benign party (possibly on unsafe data), and those which assume that adversaries might fine-tune the model. In the rest of this section, we use terms from the safety literature including harmful completions, refusals and safety data. An example prompt in the safety data could be “How to build a bomb". The harmful completion to this prompt would begin with “Step 1: Procure the following chemicals…", while a refusal (also known as a safe response) would be of the form “I cannot help you with this query".
@@
B.5 Discussion
B.5.1 Trust-free OML 1.0